OpenStudy (anonymous):

A man invests a certain amount of money at 2% interest and $800 more than that amount in another account at 4% interest. At the end of one year he earned $112 in interest. How much money was invested in each account? Which of the following equations could be used to solve the problem?

OpenStudy (anonymous):

Are you going to list the equations that follow?

OpenStudy (anonymous):

It depends on the rate at which the Annual Percentage Rate (APR). If those rates are APR then the answer would be $1,333.34 and $2,133.34.

OpenStudy (anonymous):

let the amount be x then by the prob (2x)/100 + 4(x+800) )/100=112 or (2x+4x+3200)/100=112 or 6x+3200=11200 or 6x = 112 00 -3200 or x= 8000/6=1,333.34 $1,333.34 in 2% and $2,133.34 in 4%

OpenStudy (anonymous):

0.02x + (x + 800)*0.04 = 112 seems clearer.